OpenStudy (anonymous):

use the remainder theorem to find the remainder whewn 2x^3-x^2-3x+7 is divided by x+2. state whether the binomial is a factor of the polynomial a. 25;yes b. -7;no c. 33;no d. -11;yes

OpenStudy (cwrw238):

the remainder will be the value of f(2)

OpenStudy (isaiah.feynman):

Actually f(-2)

OpenStudy (anonymous):

i know but its not the answer tho

OpenStudy (cwrw238):

yes sorry f(-2)

OpenStudy (cwrw238):

it does come to one of the options

OpenStudy (anonymous):

c? d?

OpenStudy (cwrw238):

you are guessing

OpenStudy (anonymous):

well 33 divide by 3 is 11

OpenStudy (cwrw238):

-16 -4 + 6 + 7 = ?

OpenStudy (anonymous):

its -19 but its not the answer

OpenStudy (cwrw238):

not -19

OpenStudy (isaiah.feynman):

Oops sign mistake!

OpenStudy (anonymous):

=-7 lol

OpenStudy (cwrw238):

-7

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

b. \[\frac{2 x^3-x^2-3 x+7}{x+2}=\left(2 x^2-5 x+7\right)-\frac{7}{x+2} \]

OpenStudy (cwrw238):

Remainder theorem states that the remainder when the polynomial P(x) is divided by ( x -a) the remainder is f(a)